# Vision-Language Models Represent Darker-Skinned Black Individuals as More Homogeneous than Lighter-Skinned Black Individuals

### v1

@@ 3.2 VLMs Represent Women as More Homogeneous than Men
Figure 4: Standardized cosine similarity values of Black men and women generated from all four VLMs. Higher standardized cosine similarity means more homogeneity in the stories generated for that group. Error bars represent one standard error above and below the mean.
As shown in Figure 4, stories about Black women were significantly more homogeneous than those about Black men across all four models (bs = 0.63, 0.15, 0.40, and 0.24, respectively, ps < .001). Likelihood-ratio tests revealed that including gender improved model fit for all four VLMs ($\chi^{2}$(1)s $\geq$ 7.34, ps < .01). See Table S3 for summary output of the Skin Tone and Gender Interaction Models and Table S5 for likelihood-ratio test results.
3.3 Interaction between Skin Tone and Gender
Figure 5: Standardized cosine similarity values of lighter and darker skinned men and women generated from all four VLMs. Higher standardized cosine similarity means more homogeneity in the stories generated for that group. Error bars represent one standard error above and below the mean.
As shown in Figure 5, we found mixed evidence for the interaction between skin tone and gender. In GPT-4o Mini and Llama-3.2, a positive interaction effect indicated that the impact of skin tone on homogeneity was significantly greater for women than for men (bs = 0.10 and 0.21, ps < .001). However, in GPT-4 Turbo and BLIP-3, the interaction effect was not significant (bs = -0.0048 and -0.0072, ps = .038 and .028). Likelihood-ratio tests showed that including the interaction effect improved model fit for VLMs with significant positive interactions ($\chi^{2}$(1)s = 389.86 and 838.42, ps < .001) but not for those with non-significant interactions ($\chi^{2}$(1)s = 0.76 and 1.17, ps = .38 and .28). See Table S4 for summary output of the Skin Tone and Gender Interaction Models, Table S5 for likelihood-ratio test results, and Table S6 for simple slopes analysis results.
@@ 4 Discussion
4.1 Skin Tone Bias Persists in Vision-Language Models
With the vision modality of AI models, the question of whether skin tone of individuals given as input to these AI models influences their output poses a significant threat to both stereotyping and erasure. Despite past evidence that skin tone can affect performances of gender classification systems, such questions had not been explored with respect to the more recent AI models such as VLMs. In this work, we find evidence that skin tone affects homogeneity bias, a form of stereotyping found in LLMs, suggesting that VLMs process skin tone of individuals and associated semantic attributes based solely on skin tone.
Such biases related to skin tone may partly stem from the under-representation of darker-skinned individuals in training data, a documented source of performance disparities in other AI models (e.g., Buolamwini and Gebru, 2018). Given that VLMs are primarily trained on web-scraped data, which also suffers from similar under-representation issues (see Bender et al. (2021) for a detailed review), this attribution likely applies to homogeneity bias of VLMs as well.
@@ 4.2 Convergent Evidence of Gender Homogeneity Bias
Consistent with past evidence finding that women were represented as more homogeneous relative to men in an LLM (Lee et al., 2024), we found evidence of gender homogeneity bias in VLMs. This raises the possibility that bias found in AI models that have been trained on single modalities may be passed onto multimodal models, especially when the bias remains unattended. Future work would benefit from a systematic analysis of how different modalities interact to affect homogeneity bias, thereby clarifying if added modalities on top of text amplify biases and what parts of the model architecture would be most effective for targeting bias mitigation efforts in VLMs.
4.3 The Disproportionate Effect of Skin Tone on Women
In two of four VLMs—GPT-4 Turbo and Llama-3.2—the effect of skin tone on homogeneity of group representations was significantly greater for women than for men. Upon closer inspection of the Skin Tone and Gender Interaction Models, including the interaction effect to the Skin Tone models rendered the effect of skin tone either insignificant or in the opposite direction, suggesting that the main effect of skin tone in the Skin Tone models were primarily driven by the effect of skin tone within Black women. This, to some extent, is consistent with human stereotyping patterns identified by Hill (2002), where skin tone disproportionately affects women. While Buolamwini and Gebru (2018) demonstrated intersectional bias in gender classification systems, our results demonstrate that similar biases persist in Vision-Language Models (VLMs), reinforcing the need to address intersectional bias across AI systems.
@@ 5 Limitations
A limitation of this work is the potential conflation of perceived Blackness (or racial prototypicality of Black Americans) and skin tone. Face stimuli were assigned to lighter- and darker-skinned conditions based on perceived Blackness ratings from GANFD. Although we justified this approach by emphasizing that the manipulations of face stimuli primarily reflected changes in skin tone, it remains unclear whether VLMs can detect subtle changes in facial features beyond skin tone. While darker skin tone is an important feature of perceived Blackness (Stepanova and Strube, 2012; Wilkins et al., 2010), future research may benefit from manipulations that can better isolate skin tone to enable a more controlled assessment of its impact on group representations in AI models.
Another limitation is the lack of generalizability to real, non-computer-generated faces. Although GANFD images are generally perceived as authentic, with 13.15% of raters identifying the 40 faces in this study as artificial, it is unclear if the findings apply to real-world faces. The use of GANFD images was essential for isolating skin tone effects, which is challenging with real-world faces where darker skin often correlates with other varying features, introducing confounds. GANFD images provided the necessary control by keeping other facial features consistent, allowing us to focus solely on the impact of skin tone on group representations.
Finally, a limitation of this work is that the open-source models did not consistently generate stories in response to the writing prompt, instead producing detailed descriptions of the individuals in the images. While this deviation highlights a failure to follow instructions, it also underscores the robustness of our findings. The documented biases were not confined to the intended story format but were also evident in the more objective visual descriptions generated by the open-source models. This suggests that the observed biases in group representations are pervasive and not merely artifacts of specific text formats, further highlighting the broader implications of our results.
@@ 6 Conclusion
We find that Vision-Language Models (VLMs) reproduce homogeneity bias, particularly in relation to skin tone. By analyzing computer-generated images of Black American men and women with controlled skin tone variations, we found that VLMs produced more homogeneous stories for darker-skinned individuals in three of four models. Black women were consistently represented more homogeneously than Black men, with skin tone having a greater impact on women in two models. These results highlight how biases, including homogeneity bias, persist in multimodal AI systems, emphasizing the urgent need for research to address intersectional biases in AI.
@@
Gender Skin Tone Image IDs
@@
Table S2: Summary output of the Skin Tone Models across all four VLMs. A significantly positive Skin Tone term indicates that cosine similarity was notably greater for darker-skinned individuals.
Skin Tone Model
GPT-4o mini GPT-4 Turbo Llama-3.2 BLIP-3
@@
Pair ID Intercept 0.18 0.068 0.21 0.33
Residual 0.82 0.93 0.80 0.69
Observations 499,000 499,000 499,000 499,000
Log likelihood -657,971.70 -690,385.00 -651,912.20 -616,619.10
Table S3: Summary output of the Gender Models across all four VLMs. A significantly positive Gender term indicates that cosine similarity was notably greater for women than for men.
@@
GPT-4o mini GPT-4 Turbo Llama-3.2 BLIP-3
@@
Pair ID Intercept 0.083 0.068 0.17 0.32
Residual 0.82 0.93 0.80 0.69
Observations 499,000 499,000 499,000 499,000
Log likelihood -658,069.90 -690,385.00 -652,143.80 -617,052.80
Table S4: Summary output of the Skin Tone and Gender Interaction Models across all four VLMs. A significantly positive interaction effect indicates that the effect of skin tone was significantly greater for women than for men.
Skin Tone and Gender Interaction Model
GPT-4o mini GPT-4 Turbo Llama-3.2 BLIP-3
@@
Intercept -0.29 -0.17 -0.17 -0.036
(0.039) (0.032) (0.047) (0.063)
Skin Tone -0.0070 0.16∗∗∗ -0.023∗∗∗ -0.095∗∗∗
(0.0036) (0.0039) (0.0051) (0.0047)
Gender 0.58∗∗∗ 0.23∗∗∗ 0.29∗∗∗ 0.24∗∗
(0.055) (0.045) (0.067) (0.089)
Interactions 0.10∗∗∗ -0.0048 0.21∗∗∗ -0.0072
(0.0051) (0.0055) (0.0071) (0.0067)
@@
Pair ID Intercept 0.083 0.056 0.18 0.32
Residual 0.82 0.93 0.80 0.69
Observations 499,000 499,000 499,000 499,000
Log likelihood -657,739.80 -690,379.70 -651,482.60 -616,620.40
Table S5: Results of likelihood-ratio tests.
@@
Skin Tone 881.34∗∗∗ <.001
BLIP-3 Gender 7.34∗∗ .007
Interaction 1.17 .28
Table S6: The effect of skin tone within each gender group across all VLMs.
Model Gender Effect of Skin Tone 95% CI

### v2

@@ 3.2 VLMs represent women as more homogeneous than men
Figure 4: Standardized cosine similarity values of Black men and women. Higher standardized cosine similarity means more homogeneity in the stories generated for that group. Error bars represent one standard error above and below the mean.
Stories about Black women were significantly more homogeneous than those about Black men across all VLMs (bs = 0.63, 0.15, and 0.40, respectively, ps < .001; see Figure 4). Likelihood-ratio tests revealed that including gender improved model fit for all VLMs ($\chi^{2}$(1)s $\geq$ 7.34, ps < .01). See Table S3 for summary output of the Gender Models and Table S5 for likelihood-ratio test results.
3.3 Interaction between phenotypicality and gender
Figure 5: Standardized cosine similarity values of Black men and women with lower and higher phenotypicality. Higher standardized cosine similarity means more homogeneity in the stories generated for that group. Error bars represent one standard error above and below the mean.
Finally, we found mixed evidence for the interaction between phenotypicality and gender. In GPT-4o mini and Llama-3.2, we found a positive interaction effect where the effect of phenotypicality on homogeneity was significantly greater for women than for men (bs = 0.10 and 0.21, ps < .001; see Figure 5). However, in GPT-4 Turbo, the interaction effect was not significant (b = -0.0048, p = .038). Likelihood-ratio tests showed that including the interaction effect improved model fit for VLMs with significant positive interactions ($\chi^{2}$(1)s = 389.86 and 838.42, ps < .001) but not for GPT-4 Turbo ($\chi^{2}$(1) = 0.76, p = .38). See Table S4 for summary output of the Interaction Models, Table S5 for likelihood-ratio test results, and Table S6 for simple slopes analysis results.
@@ 4 Discussion
In this work, we expanded the study of bias in VLMs to within-group biases, examining how perceived racial phenotypicality affects stereotyping in VLMs. We find that VLMs generate more homogeneous content about Black individuals with higher phenotypicality compared to those with lower phenotypicality. This pattern indicates that the level of phenotypicality in visual inputs influences the diversity of content that VLMs generate, with higher phenotypicality resulting in less individualized representations. Notably, this pattern mirrors findings from social psychology research, where humans have been shown to perceive individuals with more phenotypically Black features in more stereotypical ways. Given that VLMs are primarily trained on web-scraped data, which contains human-created content reflecting existing social biases (see Bender et al. (2021) for a detailed review), these models appear to reproduce patterns of bias similar to those observed in human perception.
@@ 4.1 Convergent evidence of gender homogeneity bias
Consistent with past evidence finding that women were represented as more homogeneous relative to men in LLMs (Lee et al., 2024) and VLMs (Lee et al., 2025), we found evidence of gender homogeneity bias in VLMs. This work extends previous findings by Lee et al. (2025) through the use of computer-generated images that allow for more controlled experimentation, enabling us to isolate phenotypical features while maintaining other facial characteristics constant. Future work would benefit from systematic analysis of what parts of the model architecture would be most effective for targeting bias mitigation efforts in VLMs.
4.2 The disproportionate effect of phenotypicality on women
In two of three VLMs—GPT-4 Turbo and Llama-3.2—the effect of phenotypicality on homogeneity of group representations was significantly greater for women than for men. Upon closer inspection of the Interaction Models, including the interaction effect to the Phenotypicality Models rendered the effect of phenotypicality either insignificant or in the opposite direction, suggesting that the main effect of phenotypicality in the Phenotypicality Models were primarily driven by the effect of phenotypicality within Black women. This shares consistencies with human stereotyping patterns where phenotypicality disproportionately affects women Hill (2002). While Buolamwini and Gebru (2018) demonstrated intersectional bias in gender classification systems, our results demonstrate that similar biases persist in Vision-Language Models (VLMs), reinforcing the importance of intersectionality in the study of AI bias.
@@ 5 Limitations
While our approach provides quantitative evidence of homogeneity bias between Black individuals with differing degrees of racial phenotypicality, we acknowledge important limitations in the metric used. Although the embedding-based cosine similarity method we used is the current standard for semantic text comparison, it still functions largely as a black box. While we present examples in Table 1 to demonstrate face validity, there remains limited transparency regarding which textual features contribute to the measured similarities. Using our measure, we can’t quite determine if certain topics, such as those related to stereotypes, are more likely to emerge for Black individuals with higher phenotypicality than those with lower. Future work could explore homogeneity bias through alternative metrics examining specific linguistic features such as word overlap, syntactic structures, and topical content, though such approaches would come with their own methodological trade-offs. Nevertheless, the field would benefit from complementary measurement approaches to triangulate how AI systems manifest homogeneity bias across different demographic groups.
Another potential concern might be our use of computer-generated rather than real faces. However, this methodological choice represents a key strength of our approach. To isolate the specific effects of perceived racial phenotypicality among Black individuals, we needed to control for all other facial features that typically covary with phenotypicality in real-world faces–a control that would be nearly impossible to achieve with real faces. The GANFD images enabled precise manipulation of phenotypicality while keeping all other facial characteristics consistent, eliminating potential confounds. This allowed us to draw more definitive conclusions about how perceived racial phenotypicality influences AI-generated representations.
@@ 6 Conclusion
Our analysis demonstrates that Vision-Language Models (VLMs) exhibit homogeneity bias influenced by perceived racial phenotypicality. Using computer-generated images of Black American men and women with systematically varied phenotypicality, we found that VLMs generate more homogeneous content about individuals with higher phenotypicality compared to those with lower phenotypicality. Our findings also reveal consistent gender disparities, with Black women represented more homogeneously than Black men across all models tested. Additionally, interaction analyses in some models showed that the effect of phenotypicality on content homogeneity was more pronounced for Black women than for Black men. These results extend our understanding of AI bias beyond traditional between-group comparisons, highlighting how within-group variations in perceived racial features influence the diversity of AI-generated representations. Our work underscores the importance of intersectionality in AI bias research and the need for more nuanced approaches to mitigate homogeneity bias in multimodal AI systems.
@@
Gender phenotypicality Image IDs
@@
Appendix S2 Appendix: Model Selection
Several models we initially collected data from, such as BLIP-3 (xgen-mm-phi3-mini-instruct-r-v1 Xue et al., 2024) and Claude 3.7 Sonnet (Anthropic, 2025), were excluded from our analysis because they refused to generate stories based on facial images. BLIP-3 produced visual descriptions instead (e.g., "The woman in the image is a beautiful black woman with curly hair and dark brown eyes. She has a serious expression and is looking at the camera."), while Claude 3.7 Sonnet declined to process facial images altogether, responding with statements like: "I notice the image contains a human face. Following my guidelines, I won’t identify or create a story about a specific individual in this photo. Instead, I can offer to write a brief fictional story about a character without referencing this specific image, or I could help with another creative request that doesn’t involve identifying the person in this photograph."
@@
Table S2: Summary output of the Phenotypicality Models across all four VLMs. In this model, lower phenotypicality was set as the reference level, with a significantly positive effect of phenotypicality indicating larger cosine similarity values for Black individuals with higher phenotypicality.
Phenotypicality Model
GPT-4o mini GPT-4 Turbo Llama-3.2
@@
Pair ID Intercept 0.18 0.068 0.21
Residual 0.82 0.93 0.80
Observations 499,000 499,000 499,000
Log likelihood -657,971.70 -690,385.00 -651,912.20
Table S3: Summary output of the Gender Models across all four VLMs. In this model, men were set as the reference level, with a significantly positive gender effect indicating higher cosine similarity values for Black women compared to Black men.
@@
GPT-4o mini GPT-4 Turbo Llama-3.2
@@
Pair ID Intercept 0.083 0.068 0.17
Residual 0.82 0.93 0.80
Observations 499,000 499,000 499,000
Log likelihood -658,069.90 -690,385.00 -652,143.80
Table S4: Summary output of the Interaction Models across all four VLMs. In this model, the phenotypicality term represents the effect of phenotypicality for men (the reference gender group), the gender term represents the effect of gender for Black individuals with lower phenotypicality (the reference phenotypicality group), and the interaction term indicates how the effect of phenotypicality differs for women compared to men.
Interaction Model
GPT-4o mini GPT-4 Turbo Llama-3.2
@@
Intercept -0.29 -0.17 -0.17
(0.039) (0.032) (0.047)
phenotypicality -0.0070 0.16∗∗∗ -0.023∗∗∗
(0.0036) (0.0039) (0.0051)
Gender 0.58∗∗∗ 0.23∗∗∗ 0.29∗∗∗
(0.055) (0.045) (0.067)
Interactions 0.10∗∗∗ -0.0048 0.21∗∗∗
(0.0051) (0.0055) (0.0071)
@@
Pair ID Intercept 0.083 0.056 0.18
Residual 0.82 0.93 0.80
Observations 499,000 499,000 499,000
Log likelihood -657,739.80 -690,379.70 -651,482.60
Table S5: Results of likelihood-ratio tests. Significant chi-square statistics indicate that including the corresponding terms significantly improves model fit, suggesting these factors have meaningful effects.
@@
Table S6: The effect of phenotypicality within each gender group across all VLMs.
Model Gender Effect of phenotypicality 95% CI